OpenStudy (anonymous):

How do you put y=0.05x^2-x+1 in general form Please help i give medals !

OpenStudy (anonymous):

Please help, i really dont get this

OpenStudy (calculusxy):

I am working on it.

OpenStudy (anonymous):

thanks :)

OpenStudy (calculusxy):

I believe the answer is 0

OpenStudy (anonymous):

It has to be put in general form and its equation is Y=a(x-h)^2+k

OpenStudy (calculusxy):

Do you know the numbers to fill the variables?

OpenStudy (anonymous):

ive gotten so far: y=0.05x^2-x+1 y-1=.05x^2-x y-1+.25=.05x^2-x+.25 y-.75=.05x^2-x+.25

OpenStudy (calculusxy):

I don't know now except for 0

OpenStudy (anonymous):

I dont think thats 0 is right

OpenStudy (calculusxy):

Sorry about that.

OpenStudy (sasogeek):

\(\large y=0.05x^{2}-x+1 \) \(\large y=0.05x^{2-x}+1 \) \(\large y=0.05x^{2-x+1} \) which one is your question?

OpenStudy (anonymous):

The very first one

OpenStudy (anonymous):

y = 0.05x(x-20) + 0.01 , u want in this form?

OpenStudy (anonymous):

i dont think so, the eqaution says y=a(x-h)^2+k

OpenStudy (calculusxy):

Is that a quadratic equation? I believe.

OpenStudy (anonymous):

im not sure, it might be called that or the general form eqaution

OpenStudy (amistre64):

general form is: y = ax^2 + bx + c

OpenStudy (anonymous):

i keep getting x=.05(x-10)-4 i think imright but im not sure

OpenStudy (amistre64):

y = 0.05x^2 - x + 1 , is in general form, unless theres some nitpicking rule about integer coeffs as opposed to ratioanl coeffs

OpenStudy (anonymous):

my paper says its in standard form

OpenStudy (amistre64):

standard form appears to be:\[y=a\left(x+\frac{b}{2a}\right)^2-\left(\frac{b^2-4ac}{4a}\right)\]

OpenStudy (calculusxy):

I believe that you should find the sum that equals to 1 and the product that's equal to 0.05. However, both the numbers used will have to be same used in the 0.05 and 1. So it can be like (x+\-___)(x+\-_____)